40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3663     Entry time: Wed Oct 6 22:46:36 2010
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: SHG at PSL table 

 I put some optics to get the green SHG on the PSL table.

Now a green light successfully comes out from the doubling crystal.

Since the optical layout of the PSL table was dramatically changed, I had to re-setup the green SHG stuff. 

 

 - what I did

I put a steering mirror after the 90/10% pick off mirror, and then a half wave plate and a convex lens.

The lens is approximately on the right place.

 To get the green beam easily, temporarily I raised the current of the NPRO up to 2 A.

I connected the oven to the heat controller, set the temperature to 36.8 deg which is the set point previously used.

After putting and aligning the oven, I finally obtained the green beam.

At the end of the work I set the NPRO current back to 0.9 A and relocked the PMC.

 

- things to be done

 1. more precise mode matching

 2. optimization of the temperature 

ELOG V3.1.3-